OpenStudy (anonymous):

Find a polar equation for the curve given by the Cartesian equation 4y = x^2 .

OpenStudy (sirm3d):

\[\large y=r\sin\theta,\quad x=r\cos\theta\]

Directrix (directrix):

@sirm3d Also, x^2 + y^2 = r^2